Two cars travel simultaneously at a common point. Car A travels at a rate of 25 m/s and accelerates at 2.5 m/s2 . Car B starts from rest and accelerates at 7.5 m/s2. How long will the two car cover the same displacement?
1
Expert's answer
2011-08-17T11:08:52-0400
Let's construct the system of two equations of motion, for the car A and B respectively:
x = x0 + v0t + a1t2/2
x = x0 + a2t2/2


0 + 25t + 2.5 t2/2 = 0 + 7.5t2/2
25t = 5t2/2
t = 10 sec